How to find n-factor for disproportionation reaction??


what is the n-factor for this reaction


Br2 + 2NaOH  -> NaOBr + NaBr + H20

Grade:12

3 Answers

Suneeta Sharma
13 years ago

Dear Student

 

you can calculate n factor in disproportionation reaction by the formula 

1/N=1/n1 +1/n2

where n1=n factor for oxidation half

          n2=n factor for reduction half

 

avinash kumar
10 years ago

please prove the formula used in disproportionation reaction

Saurabh awasthi
7 years ago
First of all balance the reaction and then apply the fomula
1/nf = 1/nf(red) + 1/nf(ox)
so 1/nf = 1/2+1/10
= 5/3